A parallel combination of a 1.41 μF capacitor and a 2.99 μF capacitor is connected in series to a 4.11 μF capacitor. This three‑capacitor combination is connected to a 15.1 V battery. Find the charge of each capacitor.
Answer #1

C' = 1.41 + 2.99 = 4.4 uF

Equivalent capacitance,

Ceq = (4.4 x 4.11)/(4.4 + 4.11)

Ceq = 2.14 uF

=====

Charge on 4.11,

Q1 = 2.14 x 15.1 = 32.31 uC

Charge on 1.41

q2 = 32.31* 1.41 / ( 1.41+ 2.99) = 10.35 uC

======

q3 = q1 - q2 = 21.96 uC

Comment in case any doubt please rate my answer....